Required information [The following information applies to the questions displayed below.] Hudson Co. reports the contributio1. Compute Hudson Co.s break-even point in units. 2. Compute Hudson Co.s break-even point in sales dollars. 1. units Break-1. Assume Hudson Co. has a target pretax income of $155,000 for 2020. What amount of sales in dollars) is needed to produce tAssume the company is considering investing in a new machine that will increase its fixed costs by $37,000 per year and decreIf the company raises its selling price to $200 per unit. 1. Compute Hudson Co.s contribution margin per unit. 2. Compute HuThe marketing manager believes that increasing advertising costs by $74,000 in 2020 will increase the companys sales volume
Answer #1

1. Break even point in unit = fixed cost / contribution per unit

= 315,000 / 35

= 9,000 units

2. Break even point in sales dollar = break even in unit x selling price per unit

= 9,000 x 175

= 1,575,000

1. sale for target income = fixed cost + target income / contribution margin per unit

= 315,000 + 155,000 / 35

= 13,429 units

amount of sale = 13,429 x 175 = 2,350,075

2. Margin of safety = actual sale - break even sale / actual sale x 100

= 2,350,075 - 1,575,000 / 2,350,075 x 100

= 32.98%